Six artificial satellites circle a space station a


Six artificial satellites circle a space station a


Rank each satellite based on the net force acting on it, rank from largest to smallest ie:


3rd> 5th> 6th


For maxiumum points, thank you.
Answer #1

1st = 200*160^2/5000 = 1024

2nd = 200*160^2/2500 = 1024

3rd = 300*80^2/10000 = 192

4th = 400*80^2/2500 = 1024

5th = 200*120^2/5000 = 576

6th = 800*40^2/10000 = 128


1st = 2nd = 4th > 5th > 3rd > 6th
Answer #2

2>1>5>4>3>6
Answer #3

Net force is directed inwards in the same direction as the acceleration.

it is the centripetal force



finding centripetal , use the equation: ac= mv^2/r

1: 200(160)^2/5000= 1024 N
2: 100(160)^2/2500= 1024 N
3: 300(80)^2/10,000= 192 N
4: 400(80)^2/2500= 1024 N
5: 200(120)^2/5000= 576 N
6: 800(40)^2/10000= 128 N


So then the net force would be ranked as would rank as: 1=2=4 >5 > 3 > 6